# rLLM: Relational Table Learning with LLMs

### v1

@@ 1. Introduction
Large language models (LLMs), like ChatGPT (brown2020language, ), are driving a new wave of artificial intelligence advancements, garnering widespread attention. These models are great at at understanding and generating text by leveraging technologies such as web-scale unsupervised pretraining (zhao2023survey, ), instruction fine-tuning (wei2022chain, ), and value alignment (wolf2023fundamental, ). Their strong performance across various tasks suggests they could be key to achieving Artificial General Intelligence (bubeck2023sparks, ).
@@
As relational databases host around 73% of the world data (dbranking, ), recent years have seen a significant shift towards Relational Table Learning (RTL) (dvzeroski2010relational, ) (zahradnik2023deep, ) (fey2023relational, )  [^2] Although “tabular” is also commonly used to describe “row-format data samples in relational databases”, our experience during development has indicated that undergraduate/graduate students, engineers, and business clients generally prefer the term “table” over “tabular”. This preference might be due to their more vocal nature. To facilitate public understanding, this paper uses the term “table”. . In this paper, we introduce the rLLM (relationLLM) project, which aims to provide a platform for rapidly developing RTL-typle methods with LLMs. As shown in Figure 2, it performs two key functions: 1) decomposing state-of-the-art Graph Neural Networks (GNNs), LLMs, and Table Neural Networks (TNNs) into standardized modules, and 2) enabling the construction of novel models in a “combine, align, and co-train” manner using these decomposed modules.
To illustrate the application of rLLM, we present a simple RTL method named BRIDGE. Specifically, this method utilizes TNNs to process table data and leverages the “foreign keys” in relational tables to construct relationships between table samples, which are then analyzed using GNNs. This approach takes into account multiple tables and the relationships between them.
@@ 2.3. Model Layer
  • Combine refers to jointly using modules from different parts. For example, (perozzi2024let, ) and (chen2023label, ) use LLMs for preliminary label annotation, and then the annotated results are fed into a GCN (gcn, ) model for graph node classification. Within the rLLM framework, we can utilize the Predictor module from the LLM part to complete the annotation task and then use the GCN module from the GNN part for the following classification task.
  • Align involves aligning the input and output feature spaces for different modules. For instance, ConGraT (brannon2023congrat, ) generates node embeddings using a language model and a GNN separately, followed by aligning these embeddings within the final embedding space. Within the rLLM framework, we can first use the Enhancer module from the LLM part to generate embeddings (and potentially perform enhancement simultaneously), then generate node embeddings with the GNN module, and finally align the two types of embeddings.
@@ 3. An Illustration Method - BRIDGE
In this section, we introduce a straightforward method named the Basic RelatIonal table-Data LearninG FramEwork (BRIDGE) to illustrate how to quickly construct RTL-type methods using rLLM.
In real-world applications of relational databases, data is usually stored in multiple (two-dimensional) tables connected by foreign keys. This means relational table data includes table features (i.e., each sample is a row in a table with many columns) and non-table features (such as the graph structure formed by foreign key relationships). Therefore, we need to handle both the table data and their interrelationships.
Firstly, for table data, we need to use table neural networks to model and learn from the data. This is due to the heterogeneity of table data, where each column’s feature type and meaning can vary significantly. As shown in Figure 4, we use a Table Encoder to model the table features. As mentioned in Section 2.2.3, we can construct different Table Encoders using the TableTransform and TableConv modules from the TNN part in rLLM, to obtain the table embeddings.
Secondly, for non-table data, particularly referring to the “foreign keys” between tables, we can leverage these foreign pivotal relationships to construct associations among multiple samples, and then use a Graph Encoder to model these connections. As mentioned in Section 2.2.1, we can use the GraphTransform and GraphConv modules from the GNN part in rLLM, to construct various Graph Encoders, ultimately producing graph embeddings. Additionally, we input the results from the TNN (i.e., table embeddings) along with other forms of non-table data into the Graph Encoder for joint processing.
@@ 4.2. Included Datasets
Compared to other datasets. The RelBench dataset (fey2023relational, ), recently scraped from websites like Amazon and Stack Overflow, includes commercially practical applications such as user lifetime value prediction and user forum activity prediction. However, this practicality also introduces some additional challenges such as imbalance, high noise levels, complex tasks, and large data volumes. In contrast, SJTUTables are enhancements of classic datasets, offering simplicity and reliable data quality. In addition, its datasets focus on the most classic classification tasks in the machine learning field, providing fixed, balanced splits to facilitate standardized evaluation. Therefore, we strongly recommend designing standard RTL methods based on SJTUTables before conducting further commercial application evaluations on the RelBench dataset.
@@ A.1. Details of Figure 1
We calculated the total data distribution based on statistics and projections from Statista [23], which estimate that the global data volume will reach a staggering 181ZB by 2025. Additionally, using reports from Chopra  [24] and IBM [25], we roughly estimated the proportion of each modality and calculated their corresponding data volumes based on the projected global data total.
@@
To calculate the LLM cost, we first determined the number of tokens that different modalities of data could be converted into. First, for pure text data, we assumed the text to be in UTF-8 encoded plain English, where each character occupies one byte. According to relevant studies [26], English words are approximately 6-11 characters. Using this ratio, we calculated the number of words corresponding to the text data and then used OpenAI’s empirical formula [27] to estimate the number of tokens. Second, for multimodal data, such as images and videos, we applied different methodologies. For images, we used 256px by 256px RGB images and employed tokenization techniques similar to those used in Vision Transformer [28], in conjunction with OpenAI’s pricing strategy for estimation[29]. For videos, we referred to Google’s Gemini 1.5 technical report [30], analyzing token calculation strategies demonstrated in the AlphaGo documentary. This approach involved randomly sampling one frame per second from the video and calculating the number of tokens in the same way as for images. Last, after determining the number of tokens for each modality, we calculated the associated costs. We assume that these data are uniformly given to GPT-3.5 Turbo[4] for analysis, the overhead required to output the token is ignored, and finally, the approximate analysis cost is obtained.

### v2

@@ 1. Introduction
Large language models (LLMs), like ChatGPT (brown2020language), are driving a new wave of artificial intelligence advancements, garnering widespread attention. These models are great at at understanding and generating text by leveraging technologies such as web-scale unsupervised pretraining (zhao2023survey), instruction fine-tuning (wei2022chain), and value alignment (wolf2023fundamental). Their strong performance across various tasks suggests they could be key to achieving Artificial General Intelligence (bubeck2023sparks).
@@
As relational databases host around 73% of the world data (dbranking), recent years have seen a significant shift towards Relational Table Learning (RTL) (dvzeroski2010relational) (zahradnik2023deep) (fey2023relational)  [^2] Although “tabular” is also commonly used to describe “row-format data samples in relational databases”, our experience during development has indicated that undergraduate/graduate students, engineers, and business clients generally prefer the term “table” over “tabular”. This preference might be due to their more vocal nature. To facilitate public understanding, this paper uses the term “table”. . In this paper, we introduce the rLLM (relationLLM) project, which aims to provide a platform for rapidly developing RTL-typle methods with LLMs. As shown in Figure 2, it performs two key functions: 1) decomposing state-of-the-art Graph Neural Networks (GNNs), LLMs, and Table Neural Networks (TNNs) into standardized modules, and 2) enabling the construction of novel models in a “combine, align, and co-train” manner using these decomposed modules.
To illustrate the application of rLLM, we present a simple illustrate RTL method named BRIDGE. The core idea is to streamline the workflow of RDL (fey2023relational) through two key simplifications: (1) considering only a single relational table connected to the target table, and (2) preprocessing all other tables into dense embeddings. Specifically, BRIDGE employs TNNs to process the target table data and leverages the foreign-key links in the selected relational table to establish relationships among table samples, which are subsequently analyzed using GNNs. This design efficiently captures inter-table dependencies with minimal architectural complexity.
@@ 2.3. Model Layer
  • Combine refers to jointly using modules from different parts. For example, (perozzi2024let) and (chen2023label) use LLMs for preliminary label annotation, and then the annotated results are fed into a GCN (gcn) model for graph node classification. Within the rLLM framework, we can utilize the Predictor module from the LLM part to complete the annotation task and then use the GCN module from the GNN part for the following classification task.
  • Align involves aligning the input and output feature spaces for different modules. For instance, ConGraT (brannon2023congrat) generates node embeddings using a language model and a GNN separately, followed by aligning these embeddings within the final embedding space. Within the rLLM framework, we can first use the Enhancer module from the LLM part to generate embeddings (and potentially perform enhancement simultaneously), then generate node embeddings with the GNN module, and finally align the two types of embeddings.
@@ 3. An Illustration Method - BRIDGE
In this section, we introduce a straightforward method named the Basic RelatIonal table-Data LearninG FramEwork (BRIDGE) to illustrate how to quickly construct RTL-type methods using rLLM. The core idea of BRIDGE is to simplify the workflow of RDL (fey2023relational) through two key simplifications: (1) only a single relational table connected to the target table is considered, and (2) all other tables, except for the target and selected relational table, are preprocessed into dense embeddings. The details are as follows.
Specifically, in real-world applications of relational databases, data is usually stored in multiple (two-dimensional) tables connected by foreign keys. This means relational table data includes table features (i.e., each sample is a row in a table with many columns) and non-table features (such as the graph structure formed by foreign key relationships). Therefore, we need to handle both the table data and their interrelationships.
Firstly, for the target table data, we need to use table neural networks to model and learn from the data. This is due to the heterogeneity of table data, where each column’s feature type and meaning can vary significantly. As shown in Figure 4, we use a Table Encoder to model the table features. As mentioned in Section 2.2.3, we can construct different Table Encoders using the TableTransform and TableConv modules from the TNN part in rLLM, to obtain the table embeddings.
Secondly, for the selected relational table, which encodes foreign-key relationships, we construct associations among samples and model them with a Graph Encoder. As described in Section 2.2.1, we can use the GraphTransform and GraphConv modules from the GNN part in rLLM, to construct various Graph Encoders, ultimately producing graph embeddings. Additionally, target table embeddings from the TNN, along with preprocessed embeddings of other tables, are also input to the Graph Encoder for joint modeling.
@@ 4. Methods and Datasets
    • Homogeneous Methods: GCN (gcn), GAT (gat), RECT (rect), TAPE (tape), and OGC (wang2023cluster).
    • HeterogeneousMethods: HAN (han) and HGT (hu2020heterogeneous).
@@
    • Single-Table Learning: TabTransformer (tabtransformer), TabNet (tabnet), and FT-Transformer (ft-transformer).
    • Relational Table Learning: BRIDGE.
@@ 4.2. Included Datasets
Compared to other datasets. The RelBench dataset (fey2023relational), recently scraped from websites like Amazon and Stack Overflow, includes commercially practical applications such as user lifetime value prediction and user forum activity prediction. However, this practicality also introduces some additional challenges such as imbalance, high noise levels, complex tasks, and large data volumes. In contrast, SJTUTables are enhancements of classic datasets, offering simplicity and reliable data quality. In addition, its datasets focus on the most classic classification tasks in the machine learning field, providing fixed, balanced splits to facilitate standardized evaluation. Therefore, we strongly recommend designing standard RTL methods based on SJTUTables before conducting further commercial application evaluations on the RelBench dataset.
@@ A.1. Details of Figure 1
We calculated the total data distribution based on statistics and projections from Statista [statista2021worldwide], which estimate that the global data volume will reach a staggering 181ZB by 2025. Additionally, using reports from Chopra  [chopra2015big] and IBM [guide2013getting], we roughly estimated the proportion of each modality and calculated their corresponding data volumes based on the projected global data total.
@@
To calculate the LLM cost, we first determined the number of tokens that different modalities of data could be converted into. First, for pure text data, we assumed the text to be in UTF-8 encoded plain English, where each character occupies one byte. According to relevant studies [smith2012distinct], English words are approximately 6-11 characters. Using this ratio, we calculated the number of words corresponding to the text data and then used OpenAI’s empirical formula [open2023tokens] to estimate the number of tokens. Second, for multimodal data, such as images and videos, we applied different methodologies. For images, we used 256px by 256px RGB images and employed tokenization techniques similar to those used in Vision Transformer [dosovitskiy2020image], in conjunction with OpenAI’s pricing strategy for estimation[open2024pricing]. For videos, we referred to Google’s Gemini 1.5 technical report [reid2024gemini], analyzing token calculation strategies demonstrated in the AlphaGo documentary. This approach involved randomly sampling one frame per second from the video and calculating the number of tokens in the same way as for images. Last, after determining the number of tokens for each modality, we calculated the associated costs. We assume that these data are uniformly given to GPT-3.5 Turbo[4] for analysis, the overhead required to output the token is ignored, and finally, the approximate analysis cost is obtained.